# How Real Are Synthetic Therapy Conversations? Evaluating Fidelity in Prolonged Exposure Dialogues

### v3

@@ 6.1 Clinical Implications
Our findings have several important implications for clinical practice. First, synthetic PE sessions can serve as valuable training tools for novice therapists learning to identify protocol deviations, as they offer controlled examples of both adherent and non-adherent interactions without privacy concerns. Second, the consistent replication of structural elements (e.g., turn-taking, session flow) indicates that synthetic data can effectively supplement clinical training materials, particularly in settings with limited access to specialized trauma training.
However, our qualitative analysis reveals limitations that automated metrics may overlook. As shown in Figure 1, even linguistically fluent synthetic dialogues can contain subtle but clinically significant fidelity violations—including role drift, premature processing, and improper SUDS implementation—that typically escape quantitative detection. These lapses would likely go unnoticed by non-specialist reviewers, highlighting the necessity of expert-guided evaluation for AI tools in clinical contexts. Finally, our PE-specific metrics provide a systematic framework for clinicians and developers to assess AI-generated content for trauma treatment, potentially setting minimum standards for therapeutic applications of synthetic data.
@@ 6.2 Qualitative Analysis of Fidelity Lapses
  • Ex. 1: Role Drift — The therapist assumes the client’s narrative responsibility, undermining the therapeutic goal of client-driven processing.
  • Ex. 2: Failure to Maintain Focus on Exposure Targets — Avoidance behavior is permitted, a lapse not distinguishable by standard metrics.
  • Ex. 3–4: Premature Reflection and Intervention — The therapist interjects before full emotional processing, disrupting the habituation process crucial for PE efficacy.
  • Ex. 5: Improper SUDS Implementation — The dialogue maintains structural flow, but the clinical assessment validity is compromised.
@@ 8 Conclusions
This study evaluated synthetic therapeutic dialogues, showing that while they replicate structural features like turn-taking, they fall short in utterance length and conversational variability. Statistical analyses confirm these gaps, revealing that surface-level fluency can mask clinically meaningful fidelity lapses. We contribute a fidelity-aware evaluation framework tailored to Prolonged Exposure (PE) therapy, along with metrics highlighting that current generative models, despite their fluency, lack the nuance needed for high-stakes therapeutic contexts. Our findings underscore the risk of overestimating quality through non-expert or automated evaluations. For NLP, this work emphasizes the need for clinically grounded benchmarks, richer linguistic modeling, and methods like expert-in-the-loop validation and time-aware metrics to improve fidelity in domain-specific generation.
@@ 9 Ethical Concerns
All annotators involved in fidelity evaluation had prior clinical training or supervision, and data access was limited to IRB-approved investigators to ensure ethical compliance. The real-world dataset used in this study was collected under IRB-approved protocols with participant consent and cannot be shared publicly. For details on the ethical safeguards, simulation design, and usage guidelines related to the synthetic dataset, we refer readers to BN et al. (2025).
@@
While our study demonstrates the promise of synthetic dialogues in approximating real PE therapy interactions, several limitations remain. Although we measure fidelity using structural, statistical, and protocol-based metrics, our evaluation does not assess therapeutic effectiveness or downstream clinical outcomes. Synthetic data that aligns structurally with real dialogues may still fall short in supporting meaningful therapeutic engagement or behavior change, especially in high-stakes or emotionally complex scenarios. Moreover, we do not evaluate inter-rater agreement on fidelity violations, which is critical for establishing the robustness of manual annotations—this is planned for future work.
Acknowledgement
This work is supported by the National Science Foundation (NSF) under Grant No. (Details omitted for double-blind review.) Any opinions, findings, & conclusions or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of the NSF.

### v4

@@ 6.1 Clinical Implications
Our findings have several important implications for clinical practice. First, synthetic PE sessions can serve as valuable training tools for novice therapists learning to identify protocol deviations, as they offer controlled examples of both adherent and non-adherent interactions without privacy concerns. Crucially, this fidelity-focused evaluation serves as an essential ethical prerequisite. Before any AI tool for therapy can be considered for patient-facing studies, a framework like ours must be used to verify its adherence to clinical protocols and prevent the risk of clinically significant failures. Second, the consistent replication of structural elements (e.g., turn-taking, session flow) indicates that synthetic data can effectively supplement clinical training materials, particularly in settings with limited access to specialized trauma training.
However, our qualitative analysis reveals limitations that automated metrics may overlook. As shown in Figure 1, even linguistically fluent synthetic dialogues can contain subtle but clinically significant fidelity violations, such as role drift, premature processing, and improper SUDS implementation, that typically escape quantitative detection. These lapses would likely go unnoticed by non-specialist reviewers, highlighting the necessity of expert-guided evaluation for AI tools in clinical contexts. Finally, our PE-specific metrics provide a systematic framework for clinicians and developers to assess AI-generated content for trauma treatment, potentially setting minimum standards for therapeutic applications of synthetic data.
@@ 6.2 Qualitative Analysis of Fidelity Lapses
  • Fidelity Lapse (Ex. 1): Role Drift - The therapist reflects on the client’s observation instead of redirecting them to the trauma narrative, which breaks the PE protocol mid-exposure.
  • Fidelity Issue (Ex. 2): Generic Affirmation - The therapist uses an overused phrase that lacks the specific grounding necessary for an anxious, compulsive client.
  • Protocol Adherence (Ex. 3): No Issue - The client’s dialogue shows they are appropriately immersed in the trauma memory, which is the goal for imaginal exposure.
  • Fidelity Lapse (Ex. 4): Reflection During Exposure - The client prematurely shifts into cognitive analysis during the exposure, a process that belongs in post-session processing.
  • Protocol Adherence (Ex. 5): Trauma Anchoring - The therapist correctly redirects the client back to the trauma scene after a drift, which successfully maintains the integrity of the PE protocol.
@@ 8 Conclusions
This study evaluated synthetic therapeutic dialogues, showing that while they replicate structural features like turn-taking, they fall short in utterance length and conversational variability. Statistical analyses confirm these gaps, revealing that surface-level fluency can mask clinically meaningful fidelity lapses. We contribute a fidelity-aware evaluation framework tailored to Prolonged Exposure (PE) therapy, along with metrics highlighting that current generative models, despite their fluency, lack the nuance needed for high-stakes therapeutic contexts. The metrics and methods in this framework are model-agnostic and can be readily applied by the research community to benchmark and improve dialogues generated from any large language model. Our findings underscore the risk of overestimating quality through non-expert or automated evaluations. For NLP, this work emphasizes the need for clinically grounded benchmarks, richer linguistic modeling, and methods like expert-in-the-loop validation and time-aware metrics to improve fidelity in domain-specific generation.
@@ 9 Ethical Concerns
All annotators involved in fidelity evaluation had prior clinical training or supervision, and data access was limited to IRB-approved investigators to ensure ethical compliance. The real-world dataset used in this study was collected under IRB-approved protocols with participant consent and cannot be shared publicly. For details on the ethical safeguards, simulation design, and usage guidelines related to the synthetic dataset, we refer readers to BN et al. (2025b).
Acknowledgement
This work was supported by the National Science Foundation (NSF) under Grant No. 2326144. Any opinions, findings, and conclusions or recommendations expressed are those of the author(s) and do not necessarily reflect the views of the NSF.
10 Data Availability
The synthetic dataset (Thousand Voices of Trauma) used in this paper is publicly available at https://huggingface.co/datasets/yenopoya/thousand-voices-trauma under the CC BY-NC 4.0 license. For collaborating on the real-world dataset, please contact PI Dr. Sherrill.
@@ 10 Data Availability
Although our study demonstrates the promise of synthetic dialogues in approximating real PE therapy interactions, several limitations remain: (1) our analysis is presented as a deep case study focused on data generated by a single, state-of-the-art LLM; (2) future work should apply our framework to compare outputs from a wider range of models to assess the generalizability of these specific findings; (3) although we measure fidelity using structural, statistical, and protocol-based metrics, our evaluation does not assess therapeutic effectiveness or downstream clinical outcomes; (4) synthetic data that aligns structurally with real dialogues may still fall short in supporting meaningful therapeutic engagement or behavior change, especially in high-stakes or emotionally complex scenarios; (5) we do not evaluate inter-rater agreement on fidelity violations, which is critical for establishing the robustness of manual annotations, and this is planned for future work.